OpenStudy (anonymous):

the center of the circle with the equation x^2+y^2+8x-10y= -5 (CAN SOMEONE HELP ME WITH THIS PLEASE)

jimthompson5910 (jim_thompson5910):

Complete the square for both x and y x^2+y^2+8x-10y= -5 x^2+8x+y^2-10y= -5 x^2+8x+16+y^2-10y= -5+16 x^2+8x+16+y^2-10y+25= -5+16+25 (x^2+8x+16)+(y^2-10y+25)= 36 (x+4)^2 + (y-5)^2 = 36 (x-(-4))^2 + (y-5)^2 = 6^2 The equation is now in the form (x-h)^2+(y-k)^2=r^2, which is a circle with center (h,k) and has a radius of r Since h = -4 and k = 5, the center of this circle is (-4, 5)

OpenStudy (raden):

if given an equation of circle : x^2+y^2+Ax+By+C=0 h=-A/2 = -8/2 = -4 k=-B/2 = -(-10)/2 = 5